2. Prove that g(x) = ( 3x + 1 ) / x has a root on [ - 1 , 1 ]


1
Expert's answer
2022-04-12T06:47:47-0400

ANSWER

First note, that the function "g(x)=\frac{3x+1}{x}" is defined on "[-1,0) \bigcup (0,1]" . On each of this intervals, the function is continuous .If "x>0" , then "g(x)>0." On the set "[-1,0)" , the values of the function are both positive and negative. For example, "g(-1)=\frac{3\cdot (-1)+1}{-1}=2>0" , "g(-0.1)=\frac{3\cdot (-0.1)+1}{-0.1}=-7<0" . On the segment "[-1,-0.1]" the function is continuous and "g(-1)<0<g(-0.1)" . By the Intermediate Value Theorem , there exists a number "c\in(-1,-0.1)" such that "g(c)=0."

Note: the root of the equation "\frac{3x+1}{x}=0" is easy of calculate by solving the equation "3x+1=0" ."c=-\frac{1}{3}" .